OpenStudy (anonymous):

the roots of f(x)=0 are 0 and 1. what are the roots of f(2x)?

OpenStudy (anonymous):

the roots of f(2x) will be 1/2 of the roots of f(x), because: \[f(2*\frac{0}{2}) = f(0) = 0, f(2*\frac{1}{2}) = f(1) = 0\] so the roots of f(2x) = 0 and 1/2

OpenStudy (anonymous):

lol i was thinking cause 2x means graph horizontally compressed by 1/2 thats why is that same reason?

OpenStudy (anonymous):

yes, that reasoning also works. Another way is also to create the quadratic equation itself.

OpenStudy (anonymous):

umm we dont know a value can we still do it?

OpenStudy (anonymous):

like vertical stretch

OpenStudy (anonymous):

if the roots are 0 and 1, the quadratic equation is: \[x^2-x\] plugging in 2x gives: \[(2x)^2-(2x) = 4x^2-2x = 2x(2x-1)\]

OpenStudy (anonymous):

ah i c

OpenStudy (anonymous):

i prefer the way I did it up top and your reasoning to that method though. Less work lol

OpenStudy (anonymous):

lol thnkx